ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 26 Liu, Zheng, Anshan Institute of Iron and Steel Technology, P. R. China Inequalities for Some Means Follow [1], using the Gauss’ and Borchart’s algorithms to consider the tmodification [2] of the Gauss arithmetic-geometric mean and the logarithmic mean. Some inequalities between them and other means are derived.

[1] J. Sandor, On certain inequalities for means, J. Math. Anal. Appl. 199 (1996), 629–635. [2] M. K. Vamanamurthy and M. Vuorinen, Inequalities for means, J. Math. Anal. Appl. 183 (1994) 155–166.